OpenStudy (anonymous):

I have the vector field F(x,y)= e^x sin(y)i + e^x cos(y)j. It is conservative, and the potential = e^x sin(y)i + e^x sin(y)j . I would like to evaluate the line integral with parameter r(t)=ln(t)i +tj from pi

OpenStudy (anonymous):

Give me just a second to work it out and I will tell you :P

OpenStudy (anonymous):

Great, thanks!

OpenStudy (anonymous):

As far as I know when evaluating line integrals in a vector field the potential is unimportant. What you are looking for is: \[\int\limits_C \vec{F}(\vec{r}(t)) \cdot \vec{r}'(t)dt\] So plugging in your parameter in your field you get: \[\vec{F}(\vec{r}(t))=e^{\ln(t)}\sin(t)\hat{i}+e^{\ln(t)}\cos(t)\hat{j}=t \sin(t)\hat{i}+t \cos(t)\hat{j}\] Now differentiating your parameter you get: \[\vec{r}'(t)=\frac{\hat{i}}{t}+\hat{j}\] Now your dot product is: \[\sin(t)+t \cos(t)\] So your integral becomes: \[\huge\int\limits_{\pi}^{\frac{3\pi}{2}}\sin(t)+t \cos(t) dt=t \sin(t)|_{\pi}^{\frac{3\pi}{2}}\] I made it big so you could see the limits and everything. Evaluating that I get negative 3pi/2. :D

OpenStudy (anonymous):

Make sense? I'm guessing thats how you did it since we got the same answer xD

OpenStudy (anonymous):

I got that too, and as well, I checked it with the fundamental theorem of line integrals. So that would lead me to believe that it is correct. Not a good answer, a GREAT one, thanks for the help!!

OpenStudy (anonymous):

You're absolutely welcome :) If you have anymore I'll do what I can to help. I'm on a tutoring listserv at my university so I'm trying to get some experience helping people. I'm glad I could be of help to you and good luck in your course/studies :)